\begin{thebibliography}{99}
\bibitem{S} Agarwal, R., O Regan, D. \& Sahu, D. 2007. Iterative
construction of fixed points of nearly asymptotically nonexpansive mappings.
Journal of Nonlinear and Convex Analysis 8: 61-79.

\bibitem{Babu} Babu, G. V. R. \& Prasad, K. N. V. V. V. 2006. Mann iteration
converges faster than Ishikawa iteration for the class of Zamfirescu
operators. Fixed Point Theory and Applications 2006(1).

\bibitem{Vasile} Berinde, V. 2007. Iterative Approximation of Fixed Points.
Springer, Berlin.

\bibitem{Berinde} Berinde, V. 2004. Picard iteration converges faster than
the Mann iteration in the class of quasi-contractive operators. Fixed Point
Theory Appl. 2: 97-105.

\bibitem{Berinde1} Berinde, V. 2004. Comparing Krasnoselskij and Mann
iterations for Lipschitzian generalized pseudocontractive operators.
Proceed. of Int. Conf. On Fixed Point Theory, Univ. of Valencia. Yokohama
Publishers.

\bibitem{Berinde2} Berinde, V. \& Berinde, M. 2005. The fastest
Krasnoselskij iteration for approximating fixed points of strictly
pseudo-contractive mappings. Carpathian J. Math 21(1-2): 13-20.

\bibitem{SS} Chang, S. S., Cho, Y. J. \& Kim, J. K. 2003. The equivalence
between the convergence of modified Picard, modified Mann, and modified
Ishikawa iterations. Mathematical and Computer Modelling 37(9): 985-991.

\bibitem{KN} Chugh, R. \& Kumar, V. 2012. Stability of hybrid fixed point
iterative algorithms of Kirk-Noor type in normed linear space for self and
nonself operators. International Journal of Contemporary Mathematical
sciences 7: 1165-1184.

\bibitem{CR} Chugh, R., Kumar, V. \& Kumar, S. 2012. Strong Convergence of a
new three step iterative scheme in Banach spaces. American Journal of
Computational Mathematics 2: 345-357.

\bibitem{Hussain1} Hussain, N., Chugh, R., Kumar, V. \& Rafiq, A. 2012. On
the rate of convergence of Kirk-type iterative schemes. Journal of Applied
Mathematics 2012.

\bibitem{Hussain} Hussain, N., Rafiq, A., Damjanovi\'{c}, B. \& Lazovi\'{c},
R. 2011. On rate of convergence of various iterative schemes. Fixed Point
Theory and Applications 2011(1): 1-6.

\bibitem{Ishikawa} Ishikawa, S. 1974. Fixed points by a new iteration
method. Proceedings of the American Mathematical Society 44: 147-150.

\bibitem{Karahan} Karahan, I. \& Ozdemir, M. 2013. A general iterative
method for approximation of fixed points and their applications. Advances in
Fixed Point Theory 3(3).

\bibitem{Karakaya} Karakaya, V., Do\u{g}an, K., G\"{u}rsoy, F. \& Ert\"{u}%
rk, M. 2013. Fixed Point of a New Three-Step Iteration Algorithm under
Contractive-Like Operators over Normed Spaces. Abstract and Applied Analysis
2013.

\bibitem{Kirk} Kirk, W. 1971. On successive approximations for nonexpansive
mappings in Banach spaces. Glasgow Math. J. 12:6-9.

\bibitem{Krasnosel'skii} Krasnosel'skii, M.A. 1955. Two remarks on the
method of successive approximations. Uspekhi Matematicheskikh Nauk 10:
123-127.

\bibitem{Mann} Mann, W.R. 1953. Mean value methods in iteration. Proceedings
of the American Mathematical Society 4: 506-510.

\bibitem{Noor} Noor, M.A. 2000. New approximation schemes for general
variational inequalities. Journal of Mathematical Analysis and Applications
251: 217-229.

\bibitem{KM-I} Olatinwo, M. 2009. Some stability results for two hybrid
fixed point iterative algorithms in normed linear space. Matematique Vesnik
61: 247-256.

\bibitem{Olatinwo} Olatinwo, M., Owojori, O. \& Imoru, C. 2006. Some
stability results for fixed point iteration processes. Aus. J. Math. Anal.
Appl. 3: 1-7.

\bibitem{Suantai} Phuengrattana, W. \& Suantai, S. 2012. Comparison of the
Rate of Convergence of Various Iterative Methods for the Class of Weak
Contractions in Banach Spaces. Thai Journal of Mathematics 11(1): 217-226.

\bibitem{SP} Phuengrattana, W. \& Suantai, S. 2011. On the rate of
convergence of Mann, Ishikawa, Noor and SP-iterations for continuous
functions on an arbitrary interval. Journal of Computational and Applied
Mathematics 235: 3006-3014.

\bibitem{Picard} Picard, E., (1890). "M\'{e}moire sur la th\'{e}orie des 
\'{e}quations aux d\'{e}riv\'{e}es partielles et la m\'{e}thode des
approximations successives", Journal de Math\'{e}matiques pures et appliqu%
\'{e}es, 6:145-210.

\bibitem{Popescu} Popescu, O. 2007. Picard iteration converges faster than
Mann iteration for a class of quasiontractive operators. Math. Commun. 12:
195-202.

\bibitem{Rhoades} Rhoades, B. E. \& \c{S}oltuz, S. M. 2004. The equivalence
between the convergences of Ishikawa and Mann iterations for an
asymptotically nonexpansive in the intermediate sense and strongly
successively pseudocontractive maps. Journal of mathematical analysis and
applications 289(1): 266-278.

\bibitem{Multistep} Rhoades, B. \& Soltuz, S.M. 2004. The equivalence
between Mann--Ishikawa iterations and multistep iteration. Nonlinear
Analysis: Theory, Methods \& Applications 58: 219-228.

\bibitem{Rhoades1} Rhoades, B. E. 1976. Comments on two fixed point
iteration methods. J. Math. Anal. Appl. 56(2): 741-750.

\bibitem{Sahu} Sahu, D. R. 2011. Applications of the S-iteration process to
constrained minimization problems and split feasibility problems. Fixed
Point Theory 12(1): 187-204.

\bibitem{GenKrasnosel'skii} Schaefer, H. 1957. \"{U}ber die Methode
sukzessiver Approximationen. Jahresbericht der Deutschen
Mathematiker-Vereinigung 59: 131-140.

\bibitem{Data Is 2} Soltuz, S.M. \& Grosan, T. 2008. Data dependence for
Ishikawa iteration when dealing with contractive like operators. Fixed Point
Theory and Applications 2008: 242916(1-7).

\bibitem{Zxue} Sun, H. \& Xue, Z. 2013. Reconsiderations on the Equivalence
of Convergence between Mann and Ishikawa Iterations for Asymptotically
Pseudocontractive Mappings. Journal of Applied Mathematics 2013.

\bibitem{iam} Thianwan, S. 2009. Common fixed points of new iterations for
two asymptotically nonexpansive nonself-mappings in a Banach space. Journal
of Computational and Applied Mathematics 224: 688-695.

\bibitem{Weng} Weng, X. 1991. Fixed point iteration for local strictly
pseudocontractive mapping. Proc. Amer. Math. Soc. 113: 727-731.

\bibitem{ZXUE} Xue, Z. 2008. The comparison of the convergence speed between
Picard, Mann, Krasnoselskij and Ishikawa iterations in Banach spaces. Fixed
Point Theory Appl. 2008.

\bibitem{Yuan} Yuan, Q. \& Rhoades, B.E. 2008. Comments on the rate of
convergence between Mann and Ishikawa iterations applied to Zamfirescu
operators. Fixed Point Theory and Applications 2008.
\end{thebibliography}

\documentclass{amsart}
\usepackage{amsmath}
\usepackage{amsfonts}

\newtheorem{theorem}{Theorem}
\theoremstyle{plain}

\newtheorem{definition}{Definition}

\newtheorem{lemma}{Lemma}

\newtheorem{remark}{Remark}

\numberwithin{equation}{section}
\input{tcilatex}

\begin{document}
\title[Convergence Analysis for Various Iterative Schemes]{Comparison of The
Speed of Convergence Among Various Iterative Schemes}
\author{Vatan KARAKAYA}
\address{Department of Mathematical Engineering, Yildiz Technical
University, Davutpasa Campus, Esenler, 34210 \.{I}stanbul, Turkey}
\email{vkkaya@yildiz.edu.tr; vkkaya@yahoo.com}
\urladdr{http://www.yarbis.yildiz.edu.tr/vkkaya}
\author{Faik G\"{U}RSOY}
\address{Department of Mathematics, Yildiz Technical University, Davutpasa
Campus, Esenler, 34220 Istanbul, Turkey}
\email{faikgursoy02@hotmail.com; fgursoy@yildiz.edu.tr}
\urladdr{http://www.yarbis.yildiz.edu.tr/fgursoy}
\author{M\"{u}zeyyen ERT\"{U}RK}
\address{Department of Mathematics, Yildiz Technical University, Davutpasa
Campus, Esenler, 34220 Istanbul, Turkey}
\email{merturk3263@gmail.com}
\urladdr{http://www.yarbis.yildiz.edu.tr/merturk}
\subjclass[2000]{Primary 47H06, 54H25.}
\keywords{Iterative schemes, Convergence, Speed of convergence, Equivalence
of convergence, Data dependence of fixed points.}

\begin{abstract}
We show that iterative scheme due to Karahan and \"{O}zdemir (2013) can be
used to approximate fixed point of contraction mappings. Furthermore, we
prove that CR iterative scheme converges faster than the iterative scheme
due to Karahan and \"{O}zdemir (2013) for the class contraction mappings.
Finally, we prove a data dependence result for contraction mappings by
employing iterative scheme due to Karahan and \"{O}zdemir (2013).
\end{abstract}

\maketitle

\section{introduction}

Fixed point theory has been appeared as one of the most powerful and
substantial theoretical tools of mathematics. This theory has a long history
and has been studied intensively by many researchers in various aspects. For
the past 30 years or so, the study of iterative procedures for the
approximation of fixed points of various classes of operators have been
flourishing areas of research for many mathematicians. Consequently,
considerable research efforts have been devoted to introduce various
iteration methods and study its more qualitative features, for example, \cite%
{S, KN, Ishikawa, Kirk, Krasnosel'skii, Mann, Noor, KM-I, Olatinwo, SP,
Picard, Multistep, GenKrasnosel'skii, iam}.

We begin our exposition with an overview of various iterative methods.

Throughout this paper $%
\mathbb{N}
$ denotes set of all nonnegative integers including zero. Let $B$ be a
Banach space, $S$ be a subset of $B$ and $T$ be a selfmap of $S$. Let $%
\left\{ \alpha _{n}^{i}\right\} _{n=0}^{\infty }$, $i\in \left\{
1,2,3\right\} $ be real sequences in $\left[ 0,1\right] $ satisfying certain
control condition(s).

An iterative sequence $\left\{ x_{n}\right\} _{n=0}^{\infty }$ defined by%
\begin{equation}
\left\{ 
\begin{array}{c}
x_{0}\in S\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ } \\ 
x_{n+1}=Tx_{n}\text{, }n\in 
\mathbb{N}
\text{,}%
\end{array}%
\right.  \label{eqn1}
\end{equation}%
is known as Picard iteration procedure \cite{Picard}, which is commonly used
to approximate fixed point of contraction mappings satisfying%
\begin{equation}
\left\Vert Tx-Ty\right\Vert \leq \delta \left\Vert x-y\right\Vert \text{, }%
\delta \in \left( 0,1\right) \text{, for all }x,y\in B\text{.}  \label{eqn2}
\end{equation}%
The following iteration methods are called Noor \cite{Noor}, and SP \cite{SP}
iteration methods, respectively:%
\begin{equation}
\left\{ 
\begin{array}{c}
x_{0}\in S\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ \ \ \ \ \ \ \ \ \ \ } \\ 
x_{n+1}=\left( 1-\alpha _{n}^{1}\right) x_{n}+\alpha _{n}^{1}Ty_{n}\text{, \
\ \ \ \ \ \ \ \ \ } \\ 
y_{n}=\left( 1-\alpha _{n}^{2}\right) x_{n}+\alpha _{n}^{2}Tz_{n}\text{, \ \
\ \ \ \ } \\ 
z_{n}=\left( 1-\alpha _{n}^{3}\right) x_{n}+\alpha _{n}^{3}Tx_{n}\text{, }%
n\in 
\mathbb{N}
\text{,}%
\end{array}%
\right.  \label{eqn3}
\end{equation}%
and%
\begin{equation}
\left\{ 
\begin{array}{c}
x_{0}\in S\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ \ \ \ \ \ \ \ \ \ \ } \\ 
x_{n+1}=\left( 1-\alpha _{n}^{1}\right) y_{n}+\alpha _{n}^{1}Ty_{n}\text{, \
\ \ \ \ \ \ \ \ \ } \\ 
y_{n}=\left( 1-\alpha _{n}^{2}\right) z_{n}+\alpha _{n}^{2}Tz_{n}\text{, \ \
\ \ \ \ } \\ 
z_{n}=\left( 1-\alpha _{n}^{3}\right) x_{n}+\alpha _{n}^{3}Tx_{n}\text{, }%
n\in 
\mathbb{N}
\text{.}%
\end{array}%
\right.  \label{eqn4}
\end{equation}

\begin{remark}
(i) Noor iteration method (1.5) reduces to well-known Ishikawa iteration
method \cite{Ishikawa} for $\alpha _{n}^{3}=0$ and Mann iteration method 
\cite{Mann} for $\alpha _{n}^{2}=\alpha _{n}^{3}=0$. (ii) SP iteration
method (1.6) reduces to a two-step Mann iteration method \cite{iam} for $%
\alpha _{n}^{3}=0$.
\end{remark}

Agarwal et al. \cite{S} inroduced an S-iteration method as follows%
\begin{equation}
\left\{ 
\begin{array}{c}
s_{0}\in S\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ \ \ \ \ \ \ \ \ \ \ } \\ 
s_{n+1}=\left( 1-\alpha _{n}^{1}\right) Ts_{n}+\alpha _{n}^{1}Tt_{n}\text{,
\ \ \ \ \ \ \ \ \ \ } \\ 
t_{n}=\left( 1-\alpha _{n}^{2}\right) s_{n}+\alpha _{n}^{2}Ts_{n}\text{, }%
n\in 
\mathbb{N}
\text{.}%
\end{array}%
\right.  \label{eqn5}
\end{equation}%
The following iteration method is referred to as CR iteration method \cite%
{CR} 
\begin{equation}
\left\{ 
\begin{array}{c}
u_{0}\in S\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ \ \ \ \ \ \ \ \ \ \ } \\ 
u_{n+1}=\left( 1-\alpha _{n}^{1}\right) v_{n}+\alpha _{n}^{1}Tv_{n}\text{, \
\ \ \ \ \ \ \ \ \ } \\ 
v_{n}=\left( 1-\alpha _{n}^{2}\right) Tu_{n}+\alpha _{n}^{2}Ty_{n}\text{, \
\ \ \ \ \ } \\ 
y_{n}=\left( 1-\alpha _{n}^{3}\right) u_{n}+\alpha _{n}^{3}Tu_{n}\text{, }%
n\in 
\mathbb{N}
\text{.}%
\end{array}%
\right.  \label{eqn6}
\end{equation}%
Very recently, Karahan and \"{O}zdemir \cite{Karahan} introduced a new three
step iteration as follows%
\begin{equation}
\left\{ 
\begin{array}{c}
p_{0}\in S\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ \ \ \ \ \ \ \ \ \ \ } \\ 
p_{n+1}=\left( 1-\alpha _{n}^{1}\right) Tp_{n}+\alpha _{n}^{1}Tq_{n}\text{,
\ \ \ \ \ \ \ \ \ \ } \\ 
q_{n}=\left( 1-\alpha _{n}^{2}\right) Tp_{n}+\alpha _{n}^{2}Tr_{n}\text{, \
\ \ \ \ \ } \\ 
r_{n}=\left( 1-\alpha _{n}^{3}\right) p_{n}+\alpha _{n}^{3}Tp_{n}\text{, }%
n\in 
\mathbb{N}
\text{.}%
\end{array}%
\right.  \label{eqn7}
\end{equation}%
Convergence analysis of iterative methods has an important role in the study
of iterative approximation of fixed point theory. Fixed point iteration
methods may exhibit radically different behaviors for various classes of
mappings. While a particular fixed point iteration method is convergent for
an appropriate class of mappings, it may not be convergent for the others.
Due to various reasons, it is important to determine whether an iteration
method converges to fixed point of a mapping. In many cases, there can be
two or more than two iteration procedures approximating to a fixed point of
a mapping, for example, \cite{SS, Karakaya, Rhoades, Zxue}. In such cases,
the critical and important point is to compare rate of convergence of these
iterations to find out which ones converge faster to that fixed point, e.g., 
\cite{Babu, Berinde, Berinde1, Berinde2, Hussain1, Hussain, Suantai,
Popescu, Rhoades1, Sahu, ZXUE, Yuan}.

Recently, several authors introduced different type iteration methods and
they have proved that their iteration methods converges faster than Picard,
Mann and Ishikawa iteration methods, e.g., \cite{CR, Karahan, Karakaya,
Suantai, Sahu}.

In this paper, we are concerned with two recent iteration methods defined by
(1.6) and (1.7). We show that iteration method (1.7) converges to fixed
point of a contraction mapping satisfying (1.2). Also, we prove that CR
iteration method (1.6) is equivalent and faster than iteration method (1.7)
for the class of contraction mappings. Finally, we give a data dependence
result for the fixed point of contraction mappings using iteration method
(1.7).

In order to obtain our main results we need following lemmas and definitions.

\begin{definition}
\cite{Vasile} Let $\left\{ a_{n}\right\} _{n=0}^{\infty }$ and $\left\{
b_{n}\right\} _{n=0}^{\infty }$ be two sequences of real numbers with limits 
$a$ and $b$, respectively. Assume that there exists%
\begin{equation}
\underset{n\rightarrow \infty }{\lim }\frac{\left\vert a_{n}-a\right\vert }{%
\left\vert b_{n}-b\right\vert }=l\text{.}  \label{eqn8}
\end{equation}

(i) If $l=0$, the we say that $\left\{ a_{n}\right\} _{n=0}^{\infty }$
converges faster to $a$ than $\left\{ b_{n}\right\} _{n=0}^{\infty }$ to $b$.

(ii) If $0<l<\infty $, then we say that $\left\{ a_{n}\right\}
_{n=0}^{\infty }$ and $\left\{ b_{n}\right\} _{n=0}^{\infty }$ have the same
rate of convergence.
\end{definition}

\begin{definition}
\cite{Vasile} Let $T$,$\widetilde{T}:B\rightarrow B$ be two operators. We
say that $\widetilde{T}$ is an approximate operator of $T$ if for all $x\in
B $ and for a fixed $\varepsilon >0$ we have 
\begin{equation}
\left\Vert Tx-\widetilde{T}x\right\Vert \leq \varepsilon .  \label{eqn9}
\end{equation}
\end{definition}

\begin{lemma}
\cite{Weng}Let $\left\{ a_{n}\right\} _{n=0}^{\infty }$ and $\left\{ \rho
_{n}\right\} _{n=0}^{\infty }$ be nonnegative real sequences satisfying the
following inequality:%
\begin{equation}
a_{n+1}\leq \left( 1-\eta _{n}\right) a_{n}+\rho _{n}\text{,}  \label{eqn10}
\end{equation}%
where $\eta _{n}\in \left( 0,1\right) $, for all $n\geq n_{0}$, $%
\dsum\nolimits_{n=1}^{\infty }\eta _{n}=\infty $, and $\frac{\rho _{n}}{\eta
_{n}}\rightarrow 0$ as $n\rightarrow \infty $. Then $\lim_{n\rightarrow
\infty }a_{n}=0$.
\end{lemma}

\begin{lemma}
\cite{Data Is 2} Let $\left\{ a_{n}\right\} _{n=0}^{\infty }$ be a
nonnegative sequence for which one assumes there exists $n_{0}\in 
\mathbb{N}
$, such that for all $n\geq n_{0}$ one has satisfied the inequality 
\begin{equation}
a_{n+1}\leq \left( 1-\mu _{n}\right) a_{n}+\mu _{n}\eta _{n}\text{,}
\label{eqn11}
\end{equation}
where $\mu _{n}\in \left( 0,1\right) ,$ for all $n\in 
\mathbb{N}
$, $\sum\limits_{n=0}^{\infty }\mu _{n}=\infty $ and $\eta _{n}\geq 0$, $%
\forall n\in 
\mathbb{N}
$. Then the following inequality holds 
\begin{equation}
0\leq \lim \sup_{n\rightarrow \infty }a_{n}\leq \lim \sup_{n\rightarrow
\infty }\eta _{n}.  \label{eqn12}
\end{equation}
\end{lemma}

\section{Main Results}

\begin{theorem}
Let $S$ be a nonempty closed convex subset of a Banach space $B$ and $%
T:S\rightarrow S$ be a contraction map satisfying condition (1.2). Let $%
\left\{ p_{n}\right\} _{n=0}^{\infty }$ be an iterative sequence generated
by (1.7) with real sequences $\left\{ \alpha _{n}^{i}\right\} _{n=0}^{\infty
}$, $i\in \left\{ 1,2,3\right\} $ in $\left[ 0,1\right] $ satisfying $%
\sum_{k=0}^{n}\alpha _{k}^{1}=\infty $. Then $\ \left\{ p_{n}\right\}
_{n=0}^{\infty }$ converges to a unique fixed point of $T$, say $x_{\ast }$.
\end{theorem}

\begin{proof}
Picard-Banach theorem guarantees the existence and uniqueness of $x_{\ast }$%
. We will show that $p_{n}\rightarrow x_{\ast }$ as $n\rightarrow \infty $.
From (1.2) and (1.7) we have%
\begin{eqnarray}
\left\Vert r_{n}-x_{\ast }\right\Vert &=&\left\Vert \left( 1-\alpha
_{n}^{3}\right) p_{n}+\alpha _{n}^{3}Tp_{n}-\left( 1-\alpha _{n}^{3}+\alpha
_{n}^{3}\right) x_{\ast }\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{3}\right) \left\Vert p_{n}-x_{\ast }\right\Vert
+\alpha _{n}^{3}\left\Vert Tp_{n}-Tx_{\ast }\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{3}\right) \left\Vert p_{n}-x_{\ast }\right\Vert
+\alpha _{n}^{3}\delta \left\Vert p_{n}-x_{\ast }\right\Vert  \notag \\
&=&\left[ 1-\alpha _{n}^{3}\left( 1-\delta \right) \right] \left\Vert
p_{n}-x_{\ast }\right\Vert \text{,}  \label{eqn13}
\end{eqnarray}%
\begin{eqnarray}
\left\Vert q_{n}-x_{\ast }\right\Vert &\leq &\left( 1-\alpha _{n}^{2}\right)
\left\Vert Tp_{n}-Tx_{\ast }\right\Vert +\alpha _{n}^{2}\left\Vert
Tr_{n}-Tx_{\ast }\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{2}\right) \delta \left\Vert p_{n}-x_{\ast
}\right\Vert +\alpha _{n}^{2}\delta \left\Vert r_{n}-x_{\ast }\right\Vert 
\notag \\
&\leq &\left\{ \left( 1-\alpha _{n}^{2}\right) \delta +\alpha _{n}^{2}\delta 
\left[ 1-\alpha _{n}^{3}\left( 1-\delta \right) \right] \right\} \left\Vert
p_{n}-x_{\ast }\right\Vert \text{,}  \label{eqn14}
\end{eqnarray}%
and%
\begin{eqnarray}
\left\Vert p_{n+1}-x_{\ast }\right\Vert &\leq &\left( 1-\alpha
_{n}^{1}\right) \left\Vert Tp_{n}-Tx_{\ast }\right\Vert +\alpha
_{n}^{1}\left\Vert Tq_{n}-Tx_{\ast }\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \delta \left\Vert p_{n}-x_{\ast
}\right\Vert +\alpha _{n}^{1}\delta \left\Vert q_{n}-x_{\ast }\right\Vert 
\notag \\
&\leq &\left\{ \left( 1-\alpha _{n}^{1}\right) \delta \right.  \notag \\
&&\left. +\alpha _{n}^{1}\delta \left\{ \left( 1-\alpha _{n}^{2}\right)
\delta +\alpha _{n}^{2}\delta \left[ 1-\alpha _{n}^{3}\left( 1-\delta
\right) \right] \right\} \right\} \left\Vert p_{n}-x_{\ast }\right\Vert
\label{eqn15}
\end{eqnarray}%
Since $\delta \in \left( 0,1\right) $ and $\alpha _{n}^{i}\in \left[ 0,1%
\right] $, for all $n\in 
\mathbb{N}
$ and for each $i\in \left\{ 1,2,3\right\} $%
\begin{equation}
1-\alpha _{n}^{2}\left( 1-\delta \right) <1\text{,}  \label{eqn16}
\end{equation}%
\begin{equation}
1-\alpha _{n}^{3}\left( 1-\delta \right) <1\text{.}  \label{eqn17}
\end{equation}%
By using $\delta \in \left( 0,1\right) $, (2.4) and (2.5) in (2.3), we obtain%
\begin{eqnarray}
\left\Vert p_{n+1}-x_{\ast }\right\Vert &\leq &\left[ \left( 1-\alpha
_{n}^{1}\right) \delta \left\{ \left( 1-\alpha _{n}^{2}\right) \delta
+\alpha _{n}^{2}\delta \right\} +\alpha _{n}^{1}\delta \right] \left\Vert
p_{n}-x_{\ast }\right\Vert  \notag \\
&\leq &\left[ \left( 1-\alpha _{n}^{1}\right) \delta \left\{ 1-\alpha
_{n}^{2}\left( 1-\delta \right) \right\} +\alpha _{n}^{1}\delta \right]
\left\Vert p_{n}-x_{\ast }\right\Vert  \notag \\
&\leq &\left[ 1-\alpha _{n}^{1}\left( 1-\delta \right) \right] \left\Vert
p_{n}-x_{\ast }\right\Vert  \notag \\
&\leq &\cdots  \notag \\
&\leq &\prod\limits_{k=0}^{n}\left[ 1-\alpha _{k}^{1}\left( 1-\delta \right) %
\right] \left\Vert p_{0}-x_{\ast }\right\Vert \text{.}  \label{eqn18}
\end{eqnarray}%
It is well-known from the classical analysis that $1-x\leq e^{-x}$ for all $%
x\in \left[ 0,1\right] $. By considering this fact together with (2.6), we
obtain%
\begin{eqnarray}
\left\Vert p_{n+1}-x_{\ast }\right\Vert &\leq &\prod\limits_{k=0}^{n}\left[
1-\alpha _{k}^{1}\left( 1-\delta \right) \right] \left\Vert p_{0}-x_{\ast
}\right\Vert  \notag \\
&\leq &\frac{\left\Vert p_{0}-x_{\ast }\right\Vert }{e^{\left( 1-\delta
\right) \sum_{k=0}^{n}\alpha _{k}^{1}}}\text{.}  \label{eqn19}
\end{eqnarray}%
Taking the limit of both sides of inequality (2.7) yields $p_{n}\rightarrow
x_{\ast }$ as $n\rightarrow \infty $.
\end{proof}

\begin{theorem}
Let $S$, $B$ and $T$ with fixed point $x_{\ast }$ be as in Theorem 1. Let $%
\{u_{n}\}_{n=0}^{\infty }$, $\{p_{n}\}_{n=0}^{\infty }$be two iterative
sequences defined by (1.6) and (1.7) with real sequences $\left\{ \alpha
_{n}^{i}\right\} _{n=0}^{\infty }$, $i\in \left\{ 1,2,3\right\} $ in $\left[
0,1\right] $ satisfying $\sum_{k=0}^{n}\alpha _{k}^{1}=\infty $. Then the
following are equivalent:

(i) the iteration method (1.7) converges to the fixed point $x_{\ast }$ of $%
T $;

(ii) the CR iteration method (1.6) converges to the fixed point $x_{\ast }$
of $T$.
\end{theorem}

\begin{proof}
We will prove (i)$\Rightarrow $(ii), that is, if iteration method (1.7)
converges to $x_{\ast }$, then CR iteration method (1.6) does too. Now by
using iteration method (1.7), CR iteration method (1.6) and condition (1.2),
we have%
\begin{eqnarray}
\left\Vert p_{n+1}-u_{n+1}\right\Vert &=&\left\Vert \left( 1-\alpha
_{n}^{1}\right) Tp_{n}+\alpha _{n}^{1}Tq_{n}-\left( 1-\alpha _{n}^{1}\right)
v_{n}-\alpha _{n}^{1}Tv_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left\Vert Tp_{n}-v_{n}\right\Vert
+\alpha _{n}^{1}\left\Vert Tq_{n}-Tv_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left\Vert p_{n}-v_{n}\right\Vert
+\alpha _{n}^{1}\delta \left\Vert q_{n}-v_{n}\right\Vert +\left( 1-\alpha
_{n}^{1}\right) \left\Vert p_{n}-Tp_{n}\right\Vert  \notag \\
&=&\left( 1-\alpha _{n}^{1}\right) \left\Vert \left( 1-\alpha
_{n}^{2}+\alpha _{n}^{2}\right) p_{n}-\left( 1-\alpha _{n}^{2}\right)
Tu_{n}-\alpha _{n}^{2}Ty_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \left\Vert q_{n}-v_{n}\right\Vert +\left( 1-\alpha
_{n}^{1}\right) \left\Vert p_{n}-Tp_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right)
\left\Vert p_{n}-Tu_{n}\right\Vert +\left( 1-\alpha _{n}^{1}\right) \alpha
_{n}^{2}\left\Vert p_{n}-Ty_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \left\Vert q_{n}-v_{n}\right\Vert +\left( 1-\alpha
_{n}^{1}\right) \left\Vert p_{n}-Tp_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right)
\left\{ \left\Vert p_{n}-Tp_{n}\right\Vert +\left\Vert
Tp_{n}-Tu_{n}\right\Vert \right\}  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\left\{ \left\Vert
p_{n}-Tp_{n}\right\Vert +\left\Vert Tp_{n}-Ty_{n}\right\Vert \right\}  \notag
\\
&&+\alpha _{n}^{1}\delta \left\Vert q_{n}-v_{n}\right\Vert +\left( 1-\alpha
_{n}^{1}\right) \left\Vert p_{n}-Tp_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right)
\delta \left\Vert p_{n}-u_{n}\right\Vert +\left( 1-\alpha _{n}^{1}\right)
\left( 1-\alpha _{n}^{2}\right) \left\Vert p_{n}-Tp_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\left\Vert
p_{n}-Tp_{n}\right\Vert +\left( 1-\alpha _{n}^{1}\right) \alpha
_{n}^{2}\delta \left\Vert p_{n}-y_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \left\Vert q_{n}-v_{n}\right\Vert +\left( 1-\alpha
_{n}^{1}\right) \left\Vert p_{n}-Tp_{n}\right\Vert  \notag \\
&=&\left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right) \delta
\left\Vert p_{n}-u_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \left\Vert \left(
1-\alpha _{n}^{3}+\alpha _{n}^{3}\right) p_{n}-\left( 1-\alpha
_{n}^{3}\right) u_{n}-\alpha _{n}^{3}Tu_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \left\Vert q_{n}-v_{n}\right\Vert +2\left( 1-\alpha
_{n}^{1}\right) \left\Vert p_{n}-Tp_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right)
\delta \left\Vert p_{n}-u_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \left( 1-\alpha
_{n}^{3}\right) \left\Vert p_{n}-u_{n}\right\Vert +\left( 1-\alpha
_{n}^{1}\right) \alpha _{n}^{2}\delta \alpha _{n}^{3}\left\Vert
p_{n}-Tu_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \left\Vert q_{n}-v_{n}\right\Vert +2\left( 1-\alpha
_{n}^{1}\right) \left\Vert p_{n}-Tp_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right)
\delta \left\Vert p_{n}-u_{n}\right\Vert +\left( 1-\alpha _{n}^{1}\right)
\alpha _{n}^{2}\delta \left( 1-\alpha _{n}^{3}\right) \left\Vert
p_{n}-u_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \alpha
_{n}^{3}\left\Vert p_{n}-Tp_{n}\right\Vert +\left( 1-\alpha _{n}^{1}\right)
\alpha _{n}^{2}\delta \alpha _{n}^{3}\delta \left\Vert p_{n}-u_{n}\right\Vert
\notag \\
&&+\alpha _{n}^{1}\delta \left\Vert q_{n}-v_{n}\right\Vert +2\left( 1-\alpha
_{n}^{1}\right) \left\Vert p_{n}-Tp_{n}\right\Vert  \notag \\
&=&\left\{ \left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right)
\delta +\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \left[
1-\alpha _{n}^{3}\left( 1-\delta \right) \right] \right\} \left\Vert
p_{n}-u_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \left\Vert q_{n}-v_{n}\right\Vert +\left( 1-\alpha
_{n}^{1}\right) \left( 2+\alpha _{n}^{2}\delta \alpha _{n}^{3}\right)
\left\Vert p_{n}-Tp_{n}\right\Vert \text{,}  \label{eqn20}
\end{eqnarray}%
\begin{eqnarray}
\left\Vert q_{n}-v_{n}\right\Vert &=&\left\Vert \left( 1-\alpha
_{n}^{2}\right) Tp_{n}+\alpha _{n}^{2}Tr_{n}-\left( 1-\alpha _{n}^{2}\right)
Tu_{n}-\alpha _{n}^{2}Ty_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{2}\right) \delta \left\Vert
p_{n}-u_{n}\right\Vert +\alpha _{n}^{2}\delta \left\Vert
r_{n}-y_{n}\right\Vert  \notag \\
&=&\left( 1-\alpha _{n}^{2}\right) \delta \left\Vert p_{n}-u_{n}\right\Vert 
\notag \\
&&+\alpha _{n}^{2}\delta \left\Vert \left( 1-\alpha _{n}^{3}\right)
p_{n}+\alpha _{n}^{3}Tp_{n}-\left( 1-\alpha _{n}^{3}\right) u_{n}-\alpha
_{n}^{3}Tu_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{2}\right) \delta \left\Vert
p_{n}-u_{n}\right\Vert +\alpha _{n}^{2}\delta \left[ 1-\alpha _{n}^{3}\left(
1-\delta \right) \right] \left\Vert p_{n}-u_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{2}\right) \left\Vert p_{n}-u_{n}\right\Vert
+\alpha _{n}^{2}\delta \left\Vert p_{n}-u_{n}\right\Vert  \notag \\
&=&\left[ 1-\alpha _{n}^{2}\left( 1-\delta \right) \right] \left\Vert
p_{n}-u_{n}\right\Vert \text{.}  \label{eqn21}
\end{eqnarray}%
Substituting (2.9) in (2.8) 
\begin{eqnarray}
\left\Vert p_{n+1}-u_{n+1}\right\Vert &\leq &\left\{ \left( 1-\alpha
_{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right) \delta +\left( 1-\alpha
_{n}^{1}\right) \alpha _{n}^{2}\delta \left[ 1-\alpha _{n}^{3}\left(
1-\delta \right) \right] \right.  \notag \\
&&\left. +\alpha _{n}^{1}\delta \left[ 1-\alpha _{n}^{2}\left( 1-\delta
\right) \right] \right\} \left\Vert p_{n}-u_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \left( 2+\alpha _{n}^{2}\delta \alpha
_{n}^{3}\right) \left\Vert p_{n}-Tp_{n}\right\Vert \text{.}  \label{eqn22}
\end{eqnarray}%
Since $\delta \in \left[ 0,1\right) $, $\alpha _{n}^{i}\in \left[ 0,1\right] 
$ for all $n\in 
\mathbb{N}
$ and for each $i\in \left\{ 1,2,3\right\} $,%
\begin{equation}
1-\alpha _{n}^{2}\left( 1-\delta \right) <1\text{,}  \label{eqn23}
\end{equation}%
\begin{equation}
1-\alpha _{n}^{3}\left( 1-\delta \right) <1\text{.}  \label{eqn24}
\end{equation}%
By applying inequalities (2.11) and (2.12) to (2.10), we obtain%
\begin{eqnarray}
\left\Vert p_{n+1}-u_{n+1}\right\Vert &\leq &\left[ 1-\alpha _{n}^{1}\left(
1-\delta \right) \right] \left\Vert p_{n}-u_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \left( 2+\alpha _{n}^{2}\delta \alpha
_{n}^{3}\right) \left\Vert p_{n}-Tp_{n}\right\Vert \text{.}  \label{eqn25}
\end{eqnarray}%
Using the fact $x_{\ast }=Tx_{\ast }$ and triangle inequality for norms, we
derive%
\begin{eqnarray}
\left\Vert p_{n}-Tp_{n}\right\Vert &=&\left\Vert p_{n}-x_{\ast }+Tx_{\ast
}-Tp_{n}\right\Vert  \notag \\
&\leq &\left\Vert p_{n}-x_{\ast }\right\Vert +\left\Vert Tx_{\ast
}-Tp_{n}\right\Vert  \notag \\
&\leq &\left( 1+\delta \right) \left\Vert p_{n}-x_{\ast }\right\Vert \text{.}
\label{eqn26}
\end{eqnarray}%
Substituting (2.14) in (2.13)%
\begin{eqnarray}
\left\Vert p_{n+1}-u_{n+1}\right\Vert &\leq &\left[ 1-\alpha _{n}^{1}\left(
1-\delta \right) \right] \left\Vert p_{n}-u_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \left( 2+\alpha _{n}^{2}\delta \alpha
_{n}^{3}\right) \left( 1+\delta \right) \left\Vert p_{n}-x_{\ast
}\right\Vert \text{.}  \label{eqn27}
\end{eqnarray}%
Denote that%
\begin{eqnarray}
a_{n} &=&\left\Vert p_{n}-u_{n}\right\Vert \text{,}  \notag \\
\eta _{n} &=&\alpha _{n}^{1}\left( 1-\delta \right) \in \left( 0,1\right) 
\text{,}  \label{eqn28} \\
\rho _{n} &=&\left( 1-\alpha _{n}^{1}\right) \left( 2+\alpha _{n}^{2}\delta
\alpha _{n}^{3}\right) \left( 1+\delta \right) \left\Vert p_{n}-x_{\ast
}\right\Vert \text{.}  \notag
\end{eqnarray}%
Thus, an application of Lemma 1 to (2.15) yields $a_{n}=\left\Vert
p_{n}-u_{n}\right\Vert \rightarrow 0$ as $n\rightarrow \infty $. Also, since 
$\left\Vert u_{n}-x_{\ast }\right\Vert \leq \left\Vert
p_{n}-u_{n}\right\Vert +\left\Vert p_{n}-x_{\ast }\right\Vert $, we have $%
\left\Vert u_{n}-x_{\ast }\right\Vert \rightarrow 0$ as $n\rightarrow \infty 
$.

Next, we will prove (ii)$\Rightarrow $(i). \ Assume that $\left\Vert
u_{n}-x_{\ast }\right\Vert \rightarrow 0$ as $n\rightarrow \infty $. It
follows from CR iteration method (1.6), iteration method (1.7) and condition
(1.2) that%
\begin{eqnarray}
\left\Vert u_{n+1}-p_{n+1}\right\Vert &=&\left\Vert \left( 1-\alpha
_{n}^{1}\right) v_{n}+\alpha _{n}^{1}Tv_{n}-\left( 1-\alpha _{n}^{1}\right)
Tp_{n}-\alpha _{n}^{1}Tq_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left\Vert v_{n}-Tp_{n}\right\Vert
+\alpha _{n}^{1}\left\Vert Tv_{n}-Tq_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left\Vert \left( 1-\alpha
_{n}^{2}\right) Tu_{n}+\alpha _{n}^{2}Ty_{n}-Tp_{n}\right\Vert +\alpha
_{n}^{1}\delta \left\Vert v_{n}-q_{n}\right\Vert  \notag \\
&=&\left( 1-\alpha _{n}^{1}\right) \left\Vert \left( 1-\alpha
_{n}^{2}\right) Tu_{n}+\alpha _{n}^{2}Ty_{n}-\left( 1-\alpha _{n}^{2}+\alpha
_{n}^{2}\right) Tp_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \left\Vert \left( 1-\alpha _{n}^{2}\right)
Tu_{n}+\alpha _{n}^{2}Ty_{n}-\left( 1-\alpha _{n}^{2}\right) Tp_{n}-\alpha
_{n}^{2}Tr_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right)
\left\Vert Tu_{n}-Tp_{n}\right\Vert +\left( 1-\alpha _{n}^{1}\right) \alpha
_{n}^{2}\left\Vert Ty_{n}-Tp_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \left( 1-\alpha _{n}^{2}\right) \left\Vert
Tu_{n}-Tp_{n}\right\Vert +\alpha _{n}^{1}\delta \alpha _{n}^{2}\left\Vert
Ty_{n}-Tr_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right)
\delta \left\Vert u_{n}-p_{n}\right\Vert +\left( 1-\alpha _{n}^{1}\right)
\alpha _{n}^{2}\delta \left\Vert y_{n}-p_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \left( 1-\alpha _{n}^{2}\right) \delta \left\Vert
u_{n}-p_{n}\right\Vert +\alpha _{n}^{1}\delta \alpha _{n}^{2}\delta
\left\Vert y_{n}-r_{n}\right\Vert  \notag \\
&=&\left\{ \left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right)
\delta +\alpha _{n}^{1}\delta \left( 1-\alpha _{n}^{2}\right) \delta
\right\} \left\Vert u_{n}-p_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \left\Vert \left(
1-\alpha _{n}^{3}\right) u_{n}+\alpha _{n}^{3}Tu_{n}-\left( 1-\alpha
_{n}^{3}+\alpha _{n}^{3}\right) p_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \alpha _{n}^{2}\delta \left\Vert \left( 1-\alpha
_{n}^{3}\right) u_{n}+\alpha _{n}^{3}Tu_{n}-\left( 1-\alpha _{n}^{3}\right)
p_{n}-\alpha _{n}^{3}Tp_{n}\right\Vert  \notag \\
&\leq &\left\{ \left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha
_{n}^{2}\right) \delta +\alpha _{n}^{1}\delta \left( 1-\alpha
_{n}^{2}\right) \delta \right\} \left\Vert u_{n}-p_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \left( 1-\alpha
_{n}^{3}\right) \left\Vert u_{n}-p_{n}\right\Vert +\left( 1-\alpha
_{n}^{1}\right) \alpha _{n}^{2}\delta \alpha _{n}^{3}\left\Vert
Tu_{n}-p_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \alpha _{n}^{2}\delta \left( 1-\alpha
_{n}^{3}\right) \left\Vert u_{n}-p_{n}\right\Vert +\alpha _{n}^{1}\delta
\alpha _{n}^{2}\delta \alpha _{n}^{3}\left\Vert Tu_{n}-Tp_{n}\right\Vert 
\notag \\
&\leq &\left\{ \left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha
_{n}^{2}\right) \delta +\alpha _{n}^{1}\delta \left( 1-\alpha
_{n}^{2}\right) \delta \right.  \notag \\
&&\left. +\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \left(
1-\alpha _{n}^{3}\right) +\left( 1-\alpha _{n}^{1}\right) \alpha
_{n}^{2}\delta \alpha _{n}^{3}\right.  \notag \\
&&\left. +\alpha _{n}^{1}\delta \alpha _{n}^{2}\delta \left( 1-\alpha
_{n}^{3}\right) +\alpha _{n}^{1}\delta \alpha _{n}^{2}\delta \alpha
_{n}^{3}\delta \right\} \left\Vert u_{n}-p_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \alpha
_{n}^{3}\left\Vert u_{n}-Tu_{n}\right\Vert  \notag \\
&=&\left\{ \left( 1-\alpha _{n}^{1}\right) \left( 1-\alpha _{n}^{2}\right)
\delta +\alpha _{n}^{1}\delta \left( 1-\alpha _{n}^{2}\right) \delta \right.
\notag \\
&&\left. +\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \left(
1-\alpha _{n}^{3}\right) +\left( 1-\alpha _{n}^{1}\right) \alpha
_{n}^{2}\delta \alpha _{n}^{3}\right.  \notag \\
&&\left. +\alpha _{n}^{1}\delta \alpha _{n}^{2}\delta \left[ 1-\alpha
_{n}^{3}\left( 1-\delta \right) \right] \right\} \left\Vert
u_{n}-p_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \alpha
_{n}^{3}\left\Vert u_{n}-Tu_{n}\right\Vert \text{.}  \label{eqn29}
\end{eqnarray}%
or,%
\begin{eqnarray}
\left\Vert u_{n+1}-p_{n+1}\right\Vert &\leq &\left\{ \left[ 1-\alpha
_{n}^{1}\left( 1-\delta \right) \right] \left( 1-\alpha _{n}^{2}\right)
\delta +\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \right.  \notag
\\
&&\left. +\alpha _{n}^{1}\delta \alpha _{n}^{2}\delta \left[ 1-\alpha
_{n}^{3}\left( 1-\delta \right) \right] \right\} \left\Vert
u_{n}-p_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \alpha
_{n}^{3}\left\Vert u_{n}-Tu_{n}\right\Vert \text{.}  \label{eqn30}
\end{eqnarray}%
Since $\delta \in \left[ 0,1\right) $, $\alpha _{n}^{i}\in \left[ 0,1\right] 
$ for all $n\in 
\mathbb{N}
$ and for each $i\in \left\{ 1,2,3\right\} $,%
\begin{equation}
1-\alpha _{n}^{2}\left( 1-\delta \right) <1\text{,}  \label{eqn31}
\end{equation}%
\begin{equation}
1-\alpha _{n}^{3}\left( 1-\delta \right) <1\text{.}  \label{eqn32}
\end{equation}%
By use of inequalities (2.19) and (2.20) in (2.18), we get%
\begin{eqnarray}
\left\Vert u_{n+1}-p_{n+1}\right\Vert &\leq &\left[ 1-\alpha _{n}^{1}\left(
1-\delta \right) \right] \left\Vert u_{n}-p_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \alpha
_{n}^{3}\left\Vert u_{n}-Tu_{n}\right\Vert \text{.}  \label{eqn33}
\end{eqnarray}%
Using the fact $x_{\ast }=Tx_{\ast }$ and triangle inequality for norms, we
derive%
\begin{equation}
\left\Vert u_{n}-Tu_{n}\right\Vert \leq \left( 1+\delta \right) \left\Vert
u_{n}-x_{\ast }\right\Vert \text{.}  \label{eqn34}
\end{equation}%
Hence, (2.21) becomes%
\begin{eqnarray}
\left\Vert u_{n+1}-p_{n+1}\right\Vert &\leq &\left[ 1-\alpha _{n}^{1}\left(
1-\delta \right) \right] \left\Vert u_{n}-p_{n}\right\Vert  \notag \\
&&+\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \alpha
_{n}^{3}\left( 1+\delta \right) \left\Vert u_{n}-x_{\ast }\right\Vert \text{.%
}  \label{eqn35}
\end{eqnarray}%
Define%
\begin{eqnarray}
a_{n} &=&\left\Vert u_{n}-p_{n}\right\Vert \text{,}  \notag \\
\eta _{n} &=&\alpha _{n}^{1}\left( 1-\delta \right) \in \left( 0,1\right) 
\text{,}  \label{eqn36} \\
\rho _{n} &=&\left( 1-\alpha _{n}^{1}\right) \alpha _{n}^{2}\delta \alpha
_{n}^{3}\left( 1+\delta \right) \left\Vert u_{n}-x_{\ast }\right\Vert \text{.%
}  \notag
\end{eqnarray}%
Thus, an application of Lemma 1 to (2.23) yields $a_{n}=\left\Vert
u_{n}-p_{n}\right\Vert \rightarrow 0$ as $n\rightarrow \infty $. Also, since 
$\left\Vert p_{n}-x_{\ast }\right\Vert \leq \left\Vert
p_{n}-u_{n}\right\Vert +\left\Vert u_{n}-x_{\ast }\right\Vert $, we have $%
\left\Vert p_{n}-x_{\ast }\right\Vert \rightarrow 0$ as $n\rightarrow \infty 
$.
\end{proof}

\begin{theorem}
Let $S$, $B$ and $T$ with fixed point $x_{\ast }$ be as in Theorem 1. Let $%
\left\{ \alpha _{n}^{i}\right\} _{n=0}^{\infty }$, $i\in \left\{
1,2,3\right\} $ be real sequences in $\left[ 0,1\right] $ satisfying (i) $%
\alpha _{1}\leq \alpha _{n}^{1}\leq 1$, $\alpha _{2}\leq \alpha _{n}^{2}\leq
1$ and $\alpha _{3}\leq \alpha _{n}^{3}\leq 1$, for all $n\in 
\mathbb{N}
$ and for some $\alpha _{1}$, $\alpha _{2}$, $\alpha _{3}>0$. For given $%
u_{0}=p_{0}\in S$, consider iterative sequences $\left\{ u_{n}\right\}
_{n=0}^{\infty }$ and $\left\{ p_{n}\right\} _{n=0}^{\infty }$ defined by
(1.6) and (1.7), respectively. Then $\left\{ p_{n}\right\} _{n=0}^{\infty }$
converges to $x_{\ast }$ faster than $\left\{ u_{n}\right\} _{n=0}^{\infty }$
does.
\end{theorem}

\begin{proof}
The following equality was obtained in (\cite{Karahan}, Theorem 1) 
\begin{equation}
b_{n}=\left\Vert p_{0}-x_{\ast }\right\Vert \delta ^{n+1}\left[ 1-\alpha
_{1}\left( 1-\delta \left( 1-\alpha _{2}\alpha _{3}\left( 1-\delta \right)
\right) \right) \right] ^{n+1}\text{.}  \label{eqn37}
\end{equation}%
Using now (1.7) and (1.2) we have%
\begin{eqnarray}
\left\Vert u_{n+1}-x_{\ast }\right\Vert &=&\left\Vert \left( 1-\alpha
_{n}^{1}\right) v_{n}+\alpha _{n}^{1}Tv_{n}-x_{\ast }\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left\Vert v_{n}-x_{\ast }\right\Vert
+\alpha _{n}^{1}\left\Vert Tv_{n}-x_{\ast }\right\Vert  \notag \\
&\leq &\left[ \left( 1-\alpha _{n}^{1}\right) +\alpha _{n}^{1}\delta \right]
\left\Vert v_{n}-x_{\ast }\right\Vert  \notag \\
&\leq &\left[ \left( 1-\alpha _{n}^{1}\right) +\alpha _{n}^{1}\delta \right]
\left\Vert \left( 1-\alpha _{n}^{2}\right) Tu_{n}+\alpha
_{n}^{2}Ty_{n}-x_{\ast }\right\Vert  \notag \\
&\leq &\left[ \left( 1-\alpha _{n}^{1}\right) +\alpha _{n}^{1}\delta \right]
\left( 1-\alpha _{n}^{2}\right) \left\Vert Tu_{n}-x_{\ast }\right\Vert 
\notag \\
&&+\left[ \left( 1-\alpha _{n}^{1}\right) +\alpha _{n}^{1}\delta \right]
\alpha _{n}^{2}\left\Vert Ty_{n}-x_{\ast }\right\Vert  \notag \\
&\leq &\left[ \left( 1-\alpha _{n}\right) +\alpha _{n}^{1}\delta \right]
\left( 1-\alpha _{n}^{2}\right) \delta \left\Vert u_{n}-x_{\ast }\right\Vert
\notag \\
&&+\left[ \left( 1-\alpha _{n}^{1}\right) +\alpha _{n}^{1}\delta \right]
\alpha _{n}^{2}\delta \left\Vert y_{n}-x_{\ast }\right\Vert  \notag \\
&\leq &\left[ \left( 1-\alpha _{n}^{1}\right) +\alpha _{n}^{1}\delta \right]
\left( 1-\alpha _{n}^{2}\right) \delta \left\Vert u_{n}-x_{\ast }\right\Vert
\notag \\
&&+\left[ \left( 1-\alpha _{n}^{1}\right) +\alpha _{n}^{1}\delta \right]
\alpha _{n}^{2}\delta \left( 1-\alpha _{n}^{3}\right) \left\Vert
u_{n}-x_{\ast }\right\Vert  \notag \\
&&+\left[ \left( 1-\alpha _{n}^{1}\right) +\alpha _{n}^{1}\delta \right]
\alpha _{n}^{2}\delta \alpha _{n}^{3}\delta \left\Vert u_{n}-x_{\ast
}\right\Vert  \notag \\
&=&\left[ \left( 1-\alpha _{n}^{1}\right) +\alpha _{n}^{1}\delta \right]
\left\{ \left( 1-\alpha _{n}^{2}\right) \delta +\alpha _{n}^{2}\delta \left(
1-\alpha _{n}^{3}\right) +\alpha _{n}^{2}\delta \alpha _{n}^{3}\delta
\right\} \left\Vert u_{n}-x_{\ast }\right\Vert  \notag \\
&=&\left[ 1-\alpha _{n}^{1}\left( 1-\delta \right) \right] \left[ 1-\alpha
_{n}^{2}\alpha _{n}^{3}\left( 1-\delta \right) \right] \delta \left\Vert
u_{n}-x_{\ast }\right\Vert  \notag \\
&\leq &\cdots  \notag \\
&\leq &\prod\limits_{k=0}^{n}\left[ 1-\alpha _{k}^{1}\left( 1-\delta \right) %
\right] \left[ 1-\alpha _{k}^{2}\alpha _{k}^{3}\left( 1-\delta \right) %
\right] \delta \left\Vert u_{0}-x_{\ast }\right\Vert \text{.}  \label{eqn38}
\end{eqnarray}%
From assumption (i), we obtain%
\begin{equation}
\left\Vert u_{n+1}-x_{\ast }\right\Vert \leq \left\Vert u_{0}-x_{\ast
}\right\Vert \delta ^{n+1}\left[ 1-\alpha _{1}\left( 1-\delta \right) \right]
^{n+1}\left[ 1-\alpha _{2}\alpha _{3}\left( 1-\delta \right) \right] ^{n+1}%
\text{.}  \label{eqn39}
\end{equation}%
Let%
\begin{equation}
a_{n}=\left\Vert u_{0}-x_{\ast }\right\Vert \delta ^{n+1}\left[ 1-\alpha
_{1}\left( 1-\delta \right) \right] ^{n+1}\left[ 1-\alpha _{2}\alpha
_{3}\left( 1-\delta \right) \right] ^{n+1}\text{.}  \label{eqn40}
\end{equation}%
Define%
\begin{eqnarray}
\theta _{n} &=&\frac{a_{n}}{b_{n}}=\frac{\left\Vert u_{0}-x_{\ast
}\right\Vert \delta ^{n+1}\left[ 1-\alpha _{1}\left( 1-\delta \right) \right]
^{n+1}\left[ 1-\alpha _{2}\alpha _{3}\left( 1-\delta \right) \right] ^{n+1}}{%
\left\Vert p_{0}-x_{\ast }\right\Vert \delta ^{n+1}\left[ 1-\alpha
_{1}\left( 1-\delta \left( 1-\alpha _{2}\alpha _{3}\left( 1-\delta \right)
\right) \right) \right] ^{n+1}}  \notag \\
&=&\frac{\left[ 1-\alpha _{1}\left( 1-\delta \right) \right] ^{n+1}\left[
1-\alpha _{2}\alpha _{3}\left( 1-\delta \right) \right] ^{n+1}}{\left[
1-\alpha _{1}\left( 1-\delta \left( 1-\alpha _{2}\alpha _{3}\left( 1-\delta
\right) \right) \right) \right] ^{n+1}}  \label{eqn41}
\end{eqnarray}%
Since $\delta \in \left( 0,1\right) $ and $\alpha _{i}\in \left( 0,1\right) $
for each $i\in \left\{ 1,2,3\right\} $%
\begin{eqnarray}
\alpha _{1} &<&1  \notag \\
&\Rightarrow &\alpha _{1}\alpha _{2}\alpha _{3}\left( 1-\delta \right)
<\alpha _{2}\alpha _{3}\left( 1-\delta \right)  \notag \\
&\Rightarrow &\alpha _{2}\alpha _{3}\left( -1+\delta \right) +\alpha
_{1}\alpha _{2}\alpha _{3}\left( 1-\delta \right) <0  \notag \\
&\Rightarrow &-\alpha _{2}\alpha _{3}+\alpha _{2}\alpha _{3}\delta +\alpha
_{1}\alpha _{2}\alpha _{3}-\alpha _{1}\alpha _{2}\alpha _{3}\delta <0  \notag
\\
&\Rightarrow &-\alpha _{2}\alpha _{3}+\alpha _{2}\alpha _{3}\delta +\alpha
_{1}\alpha _{2}\alpha _{3}-2\alpha _{1}\alpha _{2}\alpha _{3}\delta <-\alpha
_{1}\alpha _{2}\alpha _{3}\delta  \notag \\
&\Rightarrow &\left\{ 
\begin{array}{c}
1-\alpha _{1}+\alpha _{1}\delta -\alpha _{2}\alpha _{3}+\alpha _{2}\alpha
_{3}\delta +\alpha _{1}\alpha _{2}\alpha _{3}-2\alpha _{1}\alpha _{2}\alpha
_{3}\delta +\alpha _{1}\alpha _{2}\alpha _{3}\delta ^{2} \\ 
<1-\alpha _{1}+\alpha _{1}\delta -\alpha _{1}\alpha _{2}\alpha _{3}\delta
+\alpha _{1}\alpha _{2}\alpha _{3}\delta ^{2}%
\end{array}%
\right.  \notag \\
&\Rightarrow &\left\{ 
\begin{array}{c}
1-\alpha _{1}+\alpha _{1}\delta -\alpha _{2}\alpha _{3}\left( 1-\delta
\right) +\alpha _{1}\alpha _{2}\alpha _{3}\left( 1-\delta \right) ^{2} \\ 
<1-\alpha _{1}+\alpha _{1}\delta -\alpha _{1}\alpha _{2}\alpha _{3}\delta
\left( 1-\delta \right)%
\end{array}%
\right.  \notag \\
&\Rightarrow &\left\{ 
\begin{array}{c}
1-\alpha _{1}\left( 1-\delta \right) -\left[ 1-\alpha _{1}\left( 1-\delta
\right) \right] \alpha _{2}\alpha _{3}\left( 1-\delta \right) \\ 
<1-\alpha _{1}+\alpha _{1}\delta \left( 1-\alpha _{2}\alpha _{3}\left(
1-\delta \right) \right)%
\end{array}%
\right.  \notag \\
&\Rightarrow &\left[ 1-\alpha _{1}\left( 1-\delta \right) \right] \left[
1-\alpha _{2}\alpha _{3}\left( 1-\delta \right) \right] <1-\alpha _{1}\left(
1-\delta \left( 1-\alpha _{2}\alpha _{3}\left( 1-\delta \right) \right)
\right)  \notag \\
&\Rightarrow &\frac{\left[ 1-\alpha _{1}\left( 1-\delta \right) \right] %
\left[ 1-\alpha _{2}\alpha _{3}\left( 1-\delta \right) \right] }{1-\alpha
_{1}\left( 1-\delta \left( 1-\alpha _{2}\alpha _{3}\left( 1-\delta \right)
\right) \right) }<1\text{,}  \label{eqn42}
\end{eqnarray}%
and thus, we have%
\begin{eqnarray}
\lim_{n\rightarrow \infty }\frac{\theta _{n+1}}{\theta _{n}}
&=&\lim_{n\rightarrow \infty }\frac{\frac{\left[ 1-\alpha _{1}\left(
1-\delta \right) \right] ^{n+2}\left[ 1-\alpha _{2}\alpha _{3}\left(
1-\delta \right) \right] ^{n+2}}{\left[ 1-\alpha _{1}\left( 1-\delta \left(
1-\alpha _{2}\alpha _{3}\left( 1-\delta \right) \right) \right) \right]
^{n+2}}}{\frac{\left[ 1-\alpha _{1}\left( 1-\delta \right) \right] ^{n+1}%
\left[ 1-\alpha _{2}\alpha _{3}\left( 1-\delta \right) \right] ^{n+1}}{\left[
1-\alpha _{1}\left( 1-\delta \left( 1-\alpha _{2}\alpha _{3}\left( 1-\delta
\right) \right) \right) \right] ^{n+1}}}  \notag \\
&=&\lim_{n\rightarrow \infty }\frac{\left[ 1-\alpha _{1}\left( 1-\delta
\right) \right] \left[ 1-\alpha _{2}\alpha _{3}\left( 1-\delta \right) %
\right] }{1-\alpha _{1}\left( 1-\delta \left( 1-\alpha _{2}\alpha _{3}\left(
1-\delta \right) \right) \right) }  \notag \\
&=&\frac{\left[ 1-\alpha _{1}\left( 1-\delta \right) \right] \left[ 1-\alpha
_{2}\alpha _{3}\left( 1-\delta \right) \right] }{1-\alpha _{1}\left(
1-\delta \left( 1-\alpha _{2}\alpha _{3}\left( 1-\delta \right) \right)
\right) }<1\text{.}  \label{eqn43}
\end{eqnarray}%
It thus follows from ratio test that $\sum\limits_{n=0}^{\infty }\theta
_{n}<\infty $. Hence, we have $\lim_{n\rightarrow \infty }\theta _{n}=0$
which implies that $\left\{ p_{n}\right\} _{n=0}^{\infty }$ is faster than $%
\left\{ u_{n}\right\} _{n=0}^{\infty }$.
\end{proof}

We are now able to establish the following data dependence result.

\begin{theorem}
Let $\widetilde{T}$ be an approximate operator of $T$ satisfying condition
(1.2). Let $\left\{ p_{n}\right\} _{n=0}^{\infty }$ be an iterative sequence
generated by (1.7) for $T$ and define an iterative sequence $\left\{ 
\widetilde{p}_{n}\right\} _{n=0}^{\infty }$ as follows%
\begin{equation}
\left\{ 
\begin{array}{c}
\widetilde{p}_{0}\in S\text{, \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \
\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ } \\ 
\widetilde{p}_{n+1}=\left( 1-\alpha _{n}^{1}\right) \widetilde{T}\widetilde{p%
}_{n}+\alpha _{n}^{1}\widetilde{T}\widetilde{q}_{n}\text{, \ \ \ \ \ \ \ \ \
\ } \\ 
\widetilde{q}_{n}=\left( 1-\alpha _{n}^{2}\right) \widetilde{T}\widetilde{p}%
_{n}+\alpha _{n}^{2}\widetilde{T}\widetilde{r}_{n}\text{, \ \ \ \ \ \ } \\ 
\widetilde{r}_{n}=\left( 1-\alpha _{n}^{3}\right) \widetilde{p}_{n}+\alpha
_{n}^{3}\widetilde{T}\widetilde{p}_{n}\text{, }n\in 
\mathbb{N}
\text{,}%
\end{array}%
\right.  \label{eqn44}
\end{equation}%
where $\left\{ \alpha _{n}^{i}\right\} _{n=0}^{\infty }$, $i\in \left\{
1,2,3\right\} $ be real sequences in $\left[ 0,1\right] $ satisfying (i) $%
\frac{1}{2}\leq \alpha _{n}^{1}$ for all $n\in 
\mathbb{N}
$, and (ii) $\sum\limits_{n=0}^{\infty }\alpha _{n}^{1}=\infty $. If $Tp=p$
and $\widetilde{T}\widetilde{p}=\widetilde{p}$ such that $\widetilde{p}%
_{n}\rightarrow \widetilde{p}$ as $n\rightarrow \infty $, then we have%
\begin{equation}
\left\Vert p-\widetilde{p}\right\Vert \leq \frac{5\varepsilon }{1-\delta }%
\text{,}  \label{eqn45}
\end{equation}%
where $\varepsilon >0$ is a fixed number.
\end{theorem}

\begin{proof}
It follows from (1.2), (1.7), and (2.32) that%
\begin{eqnarray}
\left\Vert r_{n}-\widetilde{r}_{n}\right\Vert &=&\left\Vert \left( 1-\alpha
_{n}^{3}\right) p_{n}+\alpha _{n}^{3}Tp_{n}-\left( 1-\alpha _{n}^{3}\right) 
\widetilde{p}_{n}-\alpha _{n}^{3}\widetilde{T}\widetilde{p}_{n}\right\Vert 
\notag \\
&\leq &\left( 1-\alpha _{n}^{3}\right) \left\Vert p_{n}-\widetilde{p}%
_{n}\right\Vert +\alpha _{n}^{3}\left\Vert Tp_{n}-\widetilde{T}\widetilde{p}%
_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{3}\right) \left\Vert p_{n}-\widetilde{p}%
_{n}\right\Vert +\alpha _{n}^{3}\left\{ \left\Vert Tp_{n}-T\widetilde{p}%
_{n}\right\Vert +\left\Vert T\widetilde{p}_{n}-\widetilde{T}\widetilde{p}%
_{n}\right\Vert \right\}  \notag \\
&\leq &\left[ 1-\alpha _{n}^{3}\left( 1-\delta \right) \right] \left\Vert
p_{n}-\widetilde{p}_{n}\right\Vert +\alpha _{n}^{3}\varepsilon \text{,}
\label{eqn46}
\end{eqnarray}%
\begin{eqnarray}
\left\Vert q_{n}-\widetilde{q}_{n}\right\Vert &=&\left\Vert \left( 1-\alpha
_{n}^{2}\right) Tp_{n}+\alpha _{n}^{2}Tr_{n}-\left( 1-\alpha _{n}^{2}\right) 
\widetilde{T}\widetilde{p}_{n}-\alpha _{n}^{2}\widetilde{T}\widetilde{r}%
_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{2}\right) \left\Vert Tp_{n}-\widetilde{T}%
\widetilde{p}_{n}\right\Vert +\alpha _{n}^{2}\left\Vert Tr_{n}-\widetilde{T}%
\widetilde{r}_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{2}\right) \left\{ \left\Vert Tp_{n}-T\widetilde{%
p}_{n}\right\Vert +\left\Vert T\widetilde{p}_{n}-\widetilde{T}\widetilde{p}%
_{n}\right\Vert \right\}  \notag \\
&&+\alpha _{n}^{2}\left\{ \left\Vert Tr_{n}-T\widetilde{r}_{n}\right\Vert
+\left\Vert T\widetilde{r}_{n}-\widetilde{T}\widetilde{r}_{n}\right\Vert
\right\}  \notag \\
&\leq &\left( 1-\alpha _{n}^{2}\right) \delta \left\Vert p_{n}-\widetilde{p}%
_{n}\right\Vert +\alpha _{n}^{2}\delta \left\Vert r_{n}-\widetilde{r}%
_{n}\right\Vert +\left( 1-\alpha _{n}^{2}\right) \varepsilon +\alpha
_{n}^{2}\varepsilon \text{,}  \label{eqn47}
\end{eqnarray}%
\begin{eqnarray}
\left\Vert p_{n+1}-\widetilde{p}_{n+1}\right\Vert &=&\left\Vert \left(
1-\alpha _{n}^{1}\right) Tp_{n}+\alpha _{n}^{1}Tq_{n}-\left( 1-\alpha
_{n}^{1}\right) \widetilde{T}\widetilde{p}_{n}-\alpha _{n}^{1}\widetilde{T}%
q_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left\Vert Tp_{n}-\widetilde{T}%
\widetilde{p}_{n}\right\Vert +\alpha _{n}^{1}\left\Vert Tq_{n}-\widetilde{T}%
\widetilde{q}_{n}\right\Vert  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left\{ \left\Vert Tp_{n}-T\widetilde{%
p}_{n}\right\Vert +\left\Vert T\widetilde{p}_{n}-\widetilde{T}\widetilde{p}%
_{n}\right\Vert \right\}  \notag \\
&&+\alpha _{n}^{1}\left\{ \left\Vert Tq_{n}-T\widetilde{q}_{n}\right\Vert
+\left\Vert T\widetilde{q}_{n}-\widetilde{T}\widetilde{q}_{n}\right\Vert
\right\}  \notag \\
&\leq &\left( 1-\alpha _{n}^{1}\right) \left\{ \delta \left\Vert p_{n}-%
\widetilde{p}_{n}\right\Vert +\varepsilon \right\} +\alpha _{n}^{1}\left\{
\delta \left\Vert q_{n}-\widetilde{q}_{n}\right\Vert +\varepsilon \right\} 
\notag \\
&=&\left( 1-\alpha _{n}^{1}\right) \delta \left\Vert p_{n}-\widetilde{p}%
_{n}\right\Vert +\alpha _{n}^{1}\delta \left\Vert q_{n}-\widetilde{q}%
_{n}\right\Vert +\left( 1-\alpha _{n}^{1}\right) \varepsilon +\alpha
_{n}^{1}\varepsilon \text{.}  \label{eqn48}
\end{eqnarray}%
Combining (2.34), (2.35), and (2.36)%
\begin{eqnarray}
\left\Vert p_{n+1}-\widetilde{p}_{n+1}\right\Vert &\leq &\left\{ \left(
1-\alpha _{n}^{1}\right) \delta +\alpha _{n}^{1}\delta \left\{ \left(
1-\alpha _{n}^{2}\right) \delta +\alpha _{n}^{2}\delta \left[ 1-\alpha
_{n}^{3}\left( 1-\delta \right) \right] \right\} \right\} \left\Vert p_{n}-%
\widetilde{p}_{n}\right\Vert  \notag \\
&&+\alpha _{n}^{1}\delta \alpha _{n}^{2}\delta \alpha _{n}^{3}\varepsilon
+\alpha _{n}^{1}\delta \left( 1-\alpha _{n}^{2}\right) \varepsilon +\alpha
_{n}^{1}\delta \alpha _{n}^{2}\varepsilon +\left( 1-\alpha _{n}^{1}\right)
\varepsilon +\alpha _{n}^{1}\varepsilon  \label{eqn49}
\end{eqnarray}%
Since $\delta \in \left( 0,1\right) $, $\alpha _{n}^{i}\in \left[ 0,1\right] 
$ for each $i\in \left\{ 1,2,3\right\} $ and for all $n\in 
\mathbb{N}
$,%
\begin{equation}
1-\alpha _{n}^{3}\left( 1-\delta \right) <1\text{,}  \label{eqn50}
\end{equation}%
\begin{equation}
1-\alpha _{n}^{2}\left( 1-\delta \right) <1\text{,}  \label{eqn51}
\end{equation}%
\begin{equation}
\alpha _{n}^{2}\alpha _{n}^{3}\delta ^{2}<1\text{,}  \label{eqn52}
\end{equation}%
\begin{equation}
\left( 1-\alpha _{n}^{2}\right) \delta <1\text{,}  \label{eqn53}
\end{equation}%
\begin{equation}
\alpha _{n}^{2}\delta <1\text{,}  \label{eqn54}
\end{equation}%
and by assumption (i) we have%
\begin{equation}
1-\alpha _{n}^{1}\leq \alpha _{n}^{1}\text{.}  \label{eqn55}
\end{equation}%
Thus, an application of inequalities (2.38), (2.39), (2.40), (2.41), (2.42)
and (2.43) to (2.37) yields%
\begin{equation}
\left\Vert p_{n+1}-\widetilde{p}_{n+1}\right\Vert \leq \left[ 1-\alpha
_{n}^{1}\left( 1-\delta \right) \right] \left\Vert p_{n}-\widetilde{p}%
_{n}\right\Vert +\alpha _{n}^{1}\left( 1-\delta \right) \frac{5\varepsilon }{%
1-\delta }\text{.}  \label{eqn56}
\end{equation}%
Let us denote%
\begin{equation}
a_{n}:=\left\Vert p_{n}-\widetilde{p}_{n}\right\Vert \text{, }\mu
_{n}:=\alpha _{n}^{1}\left( 1-\delta \right) \in \left( 0,1\right) \text{.}
\label{eqn57}
\end{equation}
It follows from Lemma 2 that%
\begin{equation}
0\leq \underset{n\rightarrow \infty }{\lim \sup }\left\Vert p_{n}-\widetilde{%
p}_{n}\right\Vert \leq \underset{n\rightarrow \infty }{\lim \sup }\frac{%
5\varepsilon }{1-\delta }\text{.}  \label{eqn58}
\end{equation}%
From Theorem 1 we know that $\lim_{n\rightarrow \infty }p_{n}=x_{\ast }$.
Thus, using this fact together with the assumption $\lim_{n\rightarrow
\infty }\widetilde{p}_{n}=u_{\ast }$ we obtain%
\begin{equation}
\left\Vert x_{\ast }-u_{\ast }\right\Vert \leq \frac{5\varepsilon }{1-\delta 
}\text{.}  \label{eqn59}
\end{equation}
\end{proof}

\end{document}